Spartan Chemicals manufactures G-1, G-2, and G-3 from a joint process. Each gas can be liquified and sold for a higher price. Data on the process are as follows:

Required:

Determine the value of the missed item. (Do not round intermediate calculations.)

Product G-1 G-2 G-3 Total
Units produced 32,200 16,500 8,600 57,300
Joint costs $62,000 $124,000
Sales value at split-off $90,600 604,000
Additional costs to liquify 42,700 $34,500 18,000 95,200
Sales value if liquified 421,000 183,000 120,000 724,000

The joint costs of $62,000 for G-1 is the portion of the total joint cost of $124,000 that had been allocated to G-1.
Answer #1
Products Sales value at split-off % of Sales value at split-off Allocation of joint costs
G1 302,000 302,000/604,000 = 50% 124,000 x 50% = 62,000
G2 211,400 211,400/604,000 = 35% 124,000 x 35% = 43,400
G3 90,600 90,600/602,000 = 15% 124,000 x 15% = 18,600
$604,000 $124,000

From the facts of the question, it is clear that joints costs have been allocated on the basis of Sales value at split-off. Out of $124,000 of the joint costs, joint costs allocated to product G1 is $62,000, which represents 50% of the joint costs. On this basis, Sales value at split-off of product G1 must be $302,000 i.e. 50% of total sales value at split-off.

Sales value at split off of product G2 = Total Sales value at split-off - Sales value at split off of product G1 - Sales value at split off of product G3

= 604,000 - 302,000 - 90,600

= $211,400

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.